### 14 Mar 2026 v1

# High-speed Imaging through Turbulence with Event-based Light Fields

Abstract
This work introduces and demonstrates the first system capable of imaging fast-moving extended non-rigid objects through strong atmospheric turbulence at high frame rate. Event cameras are a novel sensing architecture capable of estimating high-speed imagery at thousands of frames per second. However, on their own event cameras are unable to disambiguate scene motion from turbulence. In this work, we overcome this limitation using event-based light field cameras: By simultaneously capturing multiple views of a scene, event-based light field cameras and machine learning-based reconstruction algorithms are able to disambiguate motion-induced dynamics, which produce events that are strongly correlated across views, from turbulence-induced dynamics, which produce events that are weakly correlated across view. Tabletop experiments demonstrate event-based light field can overcome strong turbulence while imaging high-speed objects traveling at up to 16,000 pixels per second.
Figure 1: Method overview. We introduce an event-based light field camera to image through turbulence at high speed. The event camera avoids motion blur by detecting brightness changes at microsecond time resolution. The light field offers cross-view constraints on the scene.
1 Introduction
Long-range imaging through the atmosphere is the primary goal of terrestrial astronomy, remote sensing, surveillance, and other applications. Atmospheric turbulence represents the central impediment to this goal. Regardless of sensor resolution, aperture size, platform stability, or cost, atmospheric turbulence significantly degrades image fidelity by introducing blur, geometric distortion, and wander.
Traditionally, turbulence has been mitigated with hardware-based solutions, such as adaptive optics (AO) [hampson2021adaptive], or software-based solutions, such as lucky imaging [fried1978probabilitylucky] or, more recently, deep learning [datum, mambatm, tmt]. However, almost all existing solutions share the same fundamental constraint: they rely upon conventional image sensors which integrate light over a fixed exposure window, irrecoverably blending any scene dynamics or turbulence fluctuations that occur over a time scale shorter than the exposure time. In this work, we overcome this limitation with event cameras.
Event cameras mitigate motion blur by measuring brightness change asynchronously at microsecond timescales, making them a natural candidate for recovering high-speed dynamics. Several lines of work have used frames and events to image through turbulence [boehrer2019eventatm, boehrer2021turbulence, evturb, egtm]. However, they have relied upon restrictive rigid-body assumptions about scene motion [boehrer2019eventatm, boehrer2021turbulence] or have restricted the reconstructions to the rate of the frame-based camera [evturb, egtm], thereby giving up event-camera’s central advantage.
Event-only imaging of dynamic scenes through turbulence is nontrivial. If the scene were static and all events were due to turbulence, one may be able to exploit these events to help reconstruct the scene [he2024microsaccade]. If the turbulence was static and the scene was dynamic, one could use existing event-to-video frameworks to reconstruct a (distorted) high-speed video [e2vid, firenet, e2vid+]. However, when both the turbulence and the scene are dynamic, both will produce (largely indistinguishable) events, thereby invalidating the assumptions that existing methods rely upon.
In this work, we combine event cameras with light fields to overcome this ambiguity [qu2025eventfield]. Light fields provide multiple sub-aperture views simultaneously, each corrupted by a different turbulence realization, while scene dynamics remain largely consistent across all views. Thus, each light field view provides a constraint on the singular scene image, and inconsistent motion across views can be attributed to turbulence. We exploit this structure (implicitly) using the recurrent encoder-decoder architecture E2VID [e2vid] with stacked multi-view event inputs and train it to reconstruct clean, high-speed video through turbulence.
In summary, our contributions are:
  • The first event-only method for high-speed imaging through turbulence, breaking the frame-rate barrier of conventional cameras.
  • Demonstration that light field optics provide sufficient information for removing turbulence while preserving scene motion.
  • Real (in-lab) and simulated experiments demonstrating our method significantly outperforms single-view reconstruction.
2 Related works
2.1 Light Field Imaging
Light field imaging acquires multiple simultaneous views of a scene from distinct viewpoints, encoding both spatial and angular information of light [lightfieldsurvey2016, lightfieldsurvey2017]. Camera arrays achieve this by positioning multiple independent sensors at spatially separated viewpoints [yang2002camarray]. While conceptually straightforward, they require precise inter-sensor synchronization and their physical bulk limits deployment in constrained platforms. Single-shot alternatives using microlens arrays overcome the synchronization problem by imaging all sub-apertures onto a single sensor [adelson1992singlelenslightfield, ng2005lightfieldcamera]. However, the dense lenslet grid demands precise optical alignment and manufacturing tolerances. Kaleidoscope-based designs avoid both drawbacks [manakov2013reconfigurable, qu2025eventfield]. Planar mirrors placed at the intermediate image plane fold the optical path, routing light from distinct regions of the aperture to separate areas of the sensor and producing multiple virtual viewpoints without additional optics per channel. This construction requires no microlens fabrication and eliminates inter-camera synchronization entirely, accordingly we adopt it here.
2.2 Imaging Through Aberrations with Light Fields
Ng and Hanrahan first demonstrated that light fields could be used to correct for optical aberrations back in 2006 [ng2006lfaberration]. Light fields have since been used to image through turbulence on multiple occasions [Wu2016lfthruturbulece, loktev2011specklelfturbulence, wu2019comparisonlfturbulence, Wu2022integratedsensor]. Intuitively, by capturing multiple perspectives of the same scene at the exact same time, these cameras create valuable spatial redundancy that can be used to remove turbulence and other aberrations. These systems have relied on conventional image sensors. While a few recent works have combined light fields with event cameras to perform high-speed 3D imaging [guo2024eventlfm, qu2025eventfield], ours is the first work to use event-based light field cameras to image through turbulence.
2.3 Software-based Imaging through Turbulence
Classical turbulence mitigation algorithms often use spatial registration and temporal fusion [Mao2020tmstanley, caliskan2014tmwithopticalflow, 2008tmspline] to recover imagery. Accordingly, these methods generally depend on static references or distortion-free patches and fail in dynamic scenes or strong turbulence. Learning-based methods relax these requirements: self-supervised and test-time optimization use priors such as lucky images [fried1978probabilitylucky], blind degradation estimation [nert, Li_2021_unsupervised], or pretrained mitigation models [feng2023turbugan, cai2024temporally] and adapt without paired data, but are iterative and ill-suited to high-speed use; meanwhile physics-based simulators [turbulence_p2s] enable supervised learning with fast inference via feed-forward sequence-to-sequence [tmt] or recurrent [datum, ge2025eventmamba] networks. All of these approaches remain limited by conventional sensors’ fixed exposure, which conflates scene dynamics and turbulence at high frame rates.
Event cameras, with microsecond-level temporal resolution, are a natural candidate for resolving this bottleneck. An early work, used an event camera to identify lucky regions in conventional frames [boehrer2019eventatm, boehrer2021turbulence]. Two recent works explore this direction with machine learning: EvTurb [evturb] employs events to guide single-frame restoration, while EGTM [egtm] leverages event temporal resolution to compute adaptive fusion weights for conventional frames. In both, the final image formation depends on the integration time of the conventional sensor, leaving output constrained to standard frame rates.
A fundamental bottleneck thus persists across all existing methods: event cameras alone cannot distinguish scene-induced events from turbulence-induced events. We propose an event-only approach that leverages light field imaging to resolve this ambiguity and achieve high-speed imaging through turbulence.
2.4 Hardware-based Imaging through Turbulence
AO, wherein one measures and then physically compensates for optical aberrations using deformable mirror or spatial light modulator, is the primary hardware-based approach for imaging through turbulence [hampson2021adaptive]. The majority of AO systems rely upon wavefront sensors [watnik2018wavefront, guo2024direct, chimitt2025wavefront], which can be accelerated with event cameras [kong2020shack, ziemann2024learning, wang2024neuromorphic, grose2024convolutional]. However, with a few notable exceptions [yeminy2021guidestar, feng2023neuws, xie2024wavemo, guo2024direct, jiang2026guidestar, pellizzari2026speckle], existing AO systems rely upon guidestars—bright points in the scene that can be used for calibration—which greatly limits their flexibility. Ours is the only guidestar-free turbulence mitigation approach that can operate at kHz.
2.5 Events to Video
Reconstructing intensity frames from event streams has seen rapid evolution. Early methods [e2vid, firenet, e2vid+] employ recurrent U-Net structures with ConvLSTM or ConvGRU modules to accumulate temporal state across asynchronous events, enabling continuous video reconstruction. HyperE2VID [ercan2024hypere2vid] extends this paradigm by introducing a hypernetwork that dynamically predicts convolutional kernel sizes, improving adaptability to varying event densities. To better capture long-range temporal dependencies, ETNet [etnet] adopts transformer architectures, achieving improved reconstruction quality at the cost of increased computational overhead. More recently, EventMamba [ge2025eventmamba] leverages state-space models to achieve efficient long-range modeling with linear complexity. A fundamental assumption underlies all these methods: events arise exclusively from scene dynamics or camera motion.
Under atmospheric turbulence, events no longer arise from scene motion alone. With a single camera, both turbulence and scene dynamics introduce changes in intensity, and retraining these networks on turbulent data does not resolve this fundamental ambiguity. We overcome this limitation with a light field event camera that simultaneously captures multiple views of the scene. Events caused by scene dynamics are strongly correlated across views, while events caused by turbulence are weakly correlated, providing the cross-view structure a network can exploit to disentangle the two.
3 Methods
Figure 2: Image model for event-based light field camera under turbulence. A light field camera captures NN simultaneous sub-aperture views, each observing the same scene through an independent turbulence realization, which injects different turbulence-induced fluctuations across views while leaving the scene contribution identical.
3.1 Modeling Turbulent Event Light Fields
Our imaging model is illustrated by Figure 2. A detailed description follows. The intensity at the sensor $I(\mathbf{x},t)$ is given by the incoherent imaging integral under an anisoplanatic point spread function (PSF):
$I(\mathbf{x},t)=\int_{\Omega_{s}}|h_{\mathbf{u}}(\mathbf{x},t)|^{2}\,S(\mathbf{u},t)\,d\mathbf{u},$ (1)
where $S(\mathbf{u},t)$ is the scene radiance and $h_{\mathbf{u}}(\mathbf{x},t)$ is the complex coherent PSF which incorporates the phase errors introduced by turbulence along the optical path [goodman2005introduction]. The accumulated phase error leads to the geometric warp and spatially varying blur observed at the sensor, as dictated by $h_{\mathbf{u}}(\mathbf{x},t)$ [chan2023turbulence_textbook].
This forward model is bilinear in $S$ and $|h|^{2}$: scene motion (changes in $S$) and turbulence (changes in $h$) are multiplicatively coupled, making their individual contributions ambiguous in the observed intensity $I$. Thus, recovering both from a single observation constitutes an ill-posed blind inverse problem.
Multiple views mitigate this ambiguity. If $N$ observations share a scene $S$ through separate turbulence realizations $\{h^{(i)}\}$, the system accumulates cross-view constraints on the common $S$:
$I^{(i)}(\mathbf{x},t)=\int_{\Omega_{s}}|h^{(i)}_{\mathbf{u}}(\mathbf{x},t)|^{2}\,S(\mathbf{u},t)\,d\mathbf{u},\quad i=1,\ldots,N.$ (2)
$S$ can be assumed to be the same in each view, because at long range parallax is negligible. However, turbulence affects each view differently. Rays entering each sub-aperture traverse a spatially distinct atmospheric column, imparting a unique instantaneous phase error and associated PSF $h^{(i)}_{\mathbf{u}}(t)$.
We obtain $N$ views with a light field event camera, with the goal of high speed imaging through turbulence. Following the standard event camera model [e2vid], the intensity of each view is passed through a non-linear logarithmic threshold: an event $(\mathbf{x},t,p)\in\mathcal{E}^{(i)}$ is emitted whenever the absolute log-intensity change $\Delta L^{(i)}(\mathbf{x},t)=|\log I^{(i)}(\mathbf{x},t)-\log I^{(i)}(\mathbf{x},t{-}\Delta t)|$ exceeds a threshold $C$, with polarity $p_{i}=\operatorname{sgn}(\Delta L^{(i)})$.
Figure 3: Network architecture. Each sub-aperture event stream is converted to a voxel grid $E_{k}^{(i)}\in\mathbb{R}^{B\times H\times W}$; the NN grids are stacked along the channel dimension to form the joint input $\tilde{E}_{k}\in\mathbb{R}^{NB\times H\times W}$. A recurrent convolutional encoder-decoder processes the stacked input window by window and outputs a reconstructed frame $\hat{V}_{k}$.
Figure 4: Simulation pipeline. Turbulence is added to clean frames using a turbulence simulator based on the phase-to-space (P2S) transform [turbulence_p2s]. V2E converts turbulent frames to events [v2e]. The process is repeated 9 times to simulate a $3\times 3$ light field. The network is trained on simulated data and generalizes to tabletop experimental settings.
3.2 Turbulence Mitigation Model
While the $N$ different views provide critical cross-view constraints on the shared scene $S$, the system remains underdetermined without physical regularization on $h$. Rather than designing explicit priors for this non-convex joint estimation problem, we adopt a data-driven approach: a neural network is trained to directly approximate the inverse mapping from the $N$ event streams to the clean video $V$. By training on a large dataset of simulated turbulence [turbulence_p2s], the network implicitly learns the spatial statistics of anisoplanatic turbulence and the manifold of natural scenes, leveraging cross-view consensus to resolve the bilinear ambiguity.
We implement this by adapting E2VID [e2vid] to the multi-view setting with minimal modification, so that the performance gain can be attributed to the light field geometry rather than architectural choices. In its original single-view form, E2VID takes an event stream $\mathcal{E}$ as input and outputs a reconstructed video $\hat{V}$. E2VID represents input events as a spatio-temporal voxel grid with 5 temporal bins, whose values contain temporally weighted summed event polarity.
We extend E2VID to the multi-view setting by stacking each view’s voxel representation along the temporal bin dimension as additional channels, forming a joint input illustrated in Figure 3. E2VID processes the stacked input with a recurrent convolutional encoder-decoder, whose recurrent state accumulates temporal evidence across windows, producing the reconstructed frames $\hat{V}$.
3.2.1 Training Details
The network is trained to minimize the LPIPS loss [lpips] averaged over all reconstructed frames:
$\mathcal{L}=\frac{1}{K}\sum_{k=1}^{K}\operatorname{LPIPS}(\hat{V}_{k},V_{k}).$ (3)
E2VID is traditionally trained with the calibrated perceptual loss LPIPS instead of mean squared error (MSE) because MSE losses often produce blurry images [e2vid]. The network is trained from scratch for $200$ epochs with a batch size of $8$ using the Adam optimizer [Kingma2015Adam] with a constant learning rate of $10^{-4}$ on a single NVIDIA RTX A6000 GPU.
We follow the augmentation strategy of [e2vid, e2vid+]. Geometric augmentation applies random $112\times 112$ crops and flips with probability $0.5$. Noise augmentation adds $\mathcal{N}(0,0.4)$ to $\tilde{E}_{k}$ and injects hot-pixel noise $\mathcal{N}(0,0.1)$ at up to $0.01\%$ of pixel locations. Pause augmentation zeros input events with probability $0.05$ and maintains the paused state with probability $0.9$, training the recurrent state to preserve the output without new events. Spectral normalization [yoshida2017spectralnorm] is applied to the ConvLSTM layers to prevent artifacts from accumulating as the recurrent hidden state propagates across temporal windows.
4 Experiments
4.0.1 Simulated Dataset
A synthetic light field event dataset is constructed by combining a physics-based turbulence simulator (P2S) [turbulence_p2s] with an event simulator (V2E) [v2e], illustrated by Figure 4. Clean intensity frames are sourced from the REDS dataset [Nah_2019_CVPR_Workshops_REDS], which was recorded at 120 fps. The central region is cropped and resized to $256\times 256$, serving as ground truth. For each clip, $N=9$ independent turbulence realizations are generated by running the turbulence simulator with the same parameters but different random seeds for the P2S, producing nine sequences of degraded frames. Each degraded sequence is then passed through the event simulator to produce the corresponding sub-aperture event stream $\mathcal{E}^{(i)}$. The contrast threshold $C$ is uniformly sampled from $[0.1,0.7]$ and the refractory period is set to the default $5$ ms. The temporal window $\Delta T_{k}$ in the voxel grid is set to one frame interval.
To improve generalization across turbulence conditions, we sample turbulence parameters from the ATSyn dataset distribution and adapt P2S following the turbulence simulation procedure described in the same work [datum]. First, the pixel pitch $p$ is computed from the focal length $f$, scene width, and propagation distance, and the geometric tilt is scaled by $f/p$. Second, the blur kernel is resized to match the scale specified by the ATSyn parameters. Finally, temporal correlation is introduced by relating the random seed $\boldsymbol{n}_{t}$ at time $t$ to its predecessor via $\boldsymbol{n}_{t}=\alpha\boldsymbol{n}_{t-1}+\sqrt{1-\alpha^{2}}\,\boldsymbol{\epsilon}_{t}$, where $\alpha$ is the temporal correlation coefficient and $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
The network is trained exclusively on this simulated dataset but is evaluated on both simulated and experimental data. The dataset is split into 216 training, 27 validation, and 27 test clips, each elapsing 4 seconds. Using the turbulence severity criteria in [datum], our simulated dataset consists of 60% strong, 30% medium, and 10% weak clips. A $768\times 768$ version of the dataset was generated to compare single-view reconstructions estimated from the same number of pixels as the $3\times 3\times 256\times 256$ lightfields.
Figure 5: Tabletop experimental setup. Real turbulence is generated from temperature fluctuations caused by candles and a space heater. We placed a beam splitter behind the kaleidoscope to simultaneously capture events and reference CMOS frames.
4.0.2 Tabletop Experiments
As illustrated in Figure 5, the light field event camera comprises two optical stages in series. A Canon RF 800 mm f/11 main lens focuses the scene onto an intermediate image plane. To match the exit pupil of the main lens with the entrance pupil of the pickup system, a 100 mm pupil-matching relay lens is introduced. The pickup system forms an image from the intermediate plane onto the sensor plane. Through the reflections of a 12-inch-long kaleidoscope with a 0.495-inch aperture, similar to the setup in [qu2025eventfield], a 3 × 3 array of sub-aperture views is formed. A beam splitter behind the kaleidoscope directs light simultaneously to a Prophesee EVK4 event sensor for event capture and a Blackfly S BFS-U3-13Y3C CMOS camera for reference imagery.
Turbulence is generated in the laboratory by placing a $3\times 12$ grid of tea candles starting $2.4$ inches from the main lense and a $1500$ watt space heater $32$ inches from the same lens, as illustrated in Figure 5. All experiments are conducted with targets at a distance of $30$ feet from the camera. This is not far enough to assume targets are at optical infinity, and the kaleidoscope alignment is not perfect, so each sub-aperture view is aligned to the center with a homography estimated from a calibration target presented simultaneously to each view.
Figure 6: Light fields improve temporal consistency on experimental data. In lab, we imaged a static checkerboard and rotating dot through turbulence. The left four columns show reconstructed frames at $120$ fps. The right column visualizes $x$-$t$ and $r$-$t$ plots showing the time evolution the slice marked in yellow. The light field results in straighter lines, indicating less wander.
Table 1: Simulated quantitative results. On the REDS dataset, a light field improves frame quality in comparison to a single-view configuration.
Method Resolution PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Single-view $768\times 768$ $14.84$ $0.501$ $0.5185$
Single-view $256\times 256$ $14.84$ $0.461$ $0.4607$
Light field $3\times 3\times 256\times 256$ $\mathbf{15.40}$ $\mathbf{0.505}$ $\mathbf{0.4332}$
Figure 7: Reconstruction with light field input outperforms single-view on simulated and experimental data. Each group shows the single-view reconstruction, the light field reconstruction, and the ground truth. Simulated single-view results are at $768\times 768$ resolution, which has the same number of input pixels as the $3\times 3$ light field whose views are simulated at $256\times 256$. Experimental single-view and light field results are at $176\times 176$. The light field model recovers sharper edges with fewer artifacts when compared to the single-views of either resolution. The $x$-$t$ plots below each group show the time evolution of the slice marked in yellow. Their relative smoothness show that the light field reconstruction maintains a consistent appearance over time, while the single-view output exhibits frame-to-frame instability caused by turbulence. Videos in supplemental material more clearly illustrate turbulence reduction than images.
Figure 8: Simulated high-speed scenes show that event light fields recover sharper images than frame-based methods. Motion-blur cannot be compensated for by frame-based methods, which occurs when scene dynamics or turbulence are fast. Event-based light fields operate directly on the event stream and reconstruct the underlying scene without blur caused by integration. Motion-blur is simulated by aggregating 4 conventional frames into a single blurred frame.
5 Results
5.1 Event Light Field versus Single-View Reconstruction
To demonstrate light field optics capture sufficient information to remove turbulence while preserving scene motion, we compare the light field configuration against a single-view baseline using the same reconstruction architecture trained on the central sub-aperture only.
As illustrated in Figure 7, the model trained on light field views reconstructs edges significantly better. This is apparent along structural lines; while the single-view approach distorts them into curves, the light field model consistently maintains sharp, straight edges. This improvement is particularly evident in real-world experiments, where the single-view baseline suffers from blurred edges and dimming artifacts. Furthermore, the $x$-$t$ plot demonstrates that the light field method succeeds in producing a more temporally consistent video, effectively reducing wobble. Quantitatively, Table 1 shows that the light field model outperforms the single-view baseline across all metrics (PSNR, SSIM, and LPIPS). To ensure a fair comparison, the total pixel count is matched between the two: the single-view baseline utilizes three times the spatial resolution in both the horizontal and vertical dimensions to exactly match the aggregate pixel count of the $3\times 3$ light field.
Figure 9: Event-based light field design corrects turbulence in experimental high-speed regimes. In lab, we imaged a spinning reflective stripe (top) and a bouncing ball (bottom). A $600$ fps video is reconstructed, with samples shown four frames apart. Our method successfully reconstructs frame detail while correcting for wobble, as shown in the last column. The $\theta$-$t$ and $x$-$t$ plots show a $200$ ms interval. Videos in supplemental material more clearly illustrate turbulence reduction than images.
Additionally, two real-world experiments are presented in Figure 6 to illustrate how the light field approach reconstructs consistent videos at a traditional frame rate of 120 fps. Higher frame rate results are shown later. First, a static checkerboard is reconstructed through a turbulent medium. In the $x$-$t$ plot, one can see that the single-view baseline exhibits clear temporal jitter, whereas the light field method maintains highly stable edges over time. Next, a spinning dot is displayed on a flicker-free, LCD monitor. While the monitor displays frames discretely at 120 Hz, the motion-blur inherent to LCD monitors smooths the scene’s motion. The $r$-$t$ plot is constructed by creating a slice that follows the dot’s spin. This plot demonstrates the light field configuration effectively reduces spatial wandering caused by turbulence.
5.2 Event Light Fields Overcome Motion Blur
Figure 10: High-speed video reconstruction in experimental turbulent scenes with complex dynamics. In lab, our method recovers a water balloon bursting on top of pins at $600$ fps. The displayed images are spaced 5 frames apart. Please see the videos included in the supplemental material for more clear comparison.
We test the ability of event light fields to mitigate motion blur by simulating motion blur and comparing to MambaTM [mambatm], the state-of-the-art frame-based turbulence mitigation method. High speed imagery is simulated by averaging each group of four consecutive frames.As illustrated in Figure 8, MambaTM reconstructions still contain blur and distortions. Our method, in contrast, ingests event streams, which are free from the camera integration window, simulated from the same high-speed frames, and reconstructs video directly. The most relevant prior methods for single-view turbulence mitigation using events, EvTurb [evturb] and EGTM [egtm] have not publicly released code, precluding direct comparison. These methods use both frames and events.
5.3 Recovering High-Speed Video Through Turbulence
We performed multiple in-lab, tabletop experiments to demonstrate event based light field imaging enables high-speed long range imaging through turbulence. Unless noted otherwise, we reconstruct video at $600$ fps. As shown in Figure 9, we captured a spinning stripe and a bouncing ball in lab. Since the spinning stripe rotates uniformly, the $\theta$-$t$ plot should show straight sloped lines. The single-view configuration produces a wavy trajectory, revealing residual turbulence corruption; the light field reconstruction recovers a markedly straighter slope, consistent with the expectation. For the bouncing ball, the single-view $x$-$t$ plot shows geometric distortion and loss of surface texture, while the light field reconstruction preserves the ball’s texture along the full cross-sectional trajectory.
Figure 11: Experimental reconstruction of a Nerf dart at 12,000 fps. The light field more accurately reconstructs the straight line path of a passing dart traveling at 16,000 pixels per second. $y$-$t$ plot shows 8.3 ms of data. Images spaced 16 frames apart.
Additionally, as shown in Figure 10, we captured a water balloon bursting. The light field configuration captures the complex water dynamics without confusing them with turbulence. Finally, as shown in Figure 11, we captured a Nerf dart traveling at 16,000 pixels per second. We successfully estimate the video at 12,000 fps, with the light field reconstructing a noticeably clearer and more accurate straight line trajectory.
6 Limitations
Compared to traditional cameras, event cameras offer a limited spatial resolution ($\sim$1 MP max), which our system further reduces; we trade off spatial resolution for view diversity. One could avoid this tradeoff by using multi-camera array based light field design, but this would substantially increase cost and introduce synchronization challenges. Additionally, our system system assumes the scene is (approximately) at optical infinity and that limited parallax occurs between the views. While this is a reasonable assumption for most long-range imaging through the atmosphere applications, it is invalid for other applications (e.g., biomedicine) where one might want to image through optical aberrations up close. Finally, while our system can capture and reconstruct data at thousands of frames per second, our current video reconstruction algorithm introduces a 3.4 ms processing latency at a $176\times 176$ resolution.
7 Conclusion
We introduced and demonstrated an approach to recovering high-speed video through atmospheric turbulence using a light field event camera, which does not rely upon and is not constrained by a conventional frame-based sensor. The event camera offers high temporal resolution to resolve fast dynamics. However, on its own, it cannot discern scene dynamics from turbulence. We showed that a light field design resolves this issue. Each sub-aperture sees the same scene through a different path in the atmosphere. Then, the scene signal is consistent across views while the turbulence signature varies. Experiments with neural network based event to video reconstruction confirm that event fields mitigate turbulence better than a single-view, event fields overcome the motion-blur inherent to frame-based turbulence mitigation, simulated event fields can be used to train models that work on real-data, and event fields can reconstruct complex scene dynamics without confusing them with turbulence.